Area of a sail: The area in square meters for a triangular square sail is given by A(x)=x^2+5x+6. Find A(5) If the height of the sail is x+3 meters, then what is the length of the base of the sail?
I plugged in 5 in the x variables. A(5)= 5^2 + 5(5) +6 which will equal A(5)=56
Thus, the answer that I got was 11.2 meters. The second question is where I think I am off track. Starting with a^2+b^2=c^2 Do I plug in the 5? I was orignially using the 5 as the hypotenuse. My answers are coming out crazy. I have worked this problem for 2 hours now. I am ready to throw my computer.

Answer by scott8148(6628)

you are over thinking this

"the height of the sail is x+3 meters" __ 5+3=8

area of a triangle is 1/2 * base * height

56 = 1/2 * b * 8

b = 14


BUT--the question may be asking for the general solution

x^2+5x+6 = 1/2 * b * x+3

dividing by x+3 __ x+2 = 1/2 * b

b = 2x+4